OpenStudy (anonymous):

MEDAL!! Given f(x) = 5x − 7 and g(x) = 11x − 1, solve for (f + g)(x) and select the correct answer below. (f + g)(x) = 16x − 6 (f + g)(x) = 16x + 8 (f + g)(x) = 16x − 8 (f + g)(x) = 6x − 6

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

Please help!!

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

I'd say B or C.

OpenStudy (anonymous):

Which one makes more sense? I'm clueless

undeadknight26 (undeadknight26):

Really your just taking the two equations and combining them.

undeadknight26 (undeadknight26):

I would say C but I'm really rusty on this type of stuff.

OpenStudy (anonymous):

I was thinking the same, do you think you could help me with two more?

undeadknight26 (undeadknight26):

I could try mate.

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

If (9x − 4)(9x + 4) = ax2 − b, what is the value of a? The value of a is _____________. Numerical Answers Expected! Answer for Blank 1:

undeadknight26 (undeadknight26):

"Numerical Answers Expected!" Darn...I thought it was llama...

OpenStudy (anonymous):

Lol same, i guess they mean take the answer and put it into a numerical answer

undeadknight26 (undeadknight26):

A = 81 Taken from: http://www.wolframalpha.com/input/?i=%289x+%E2%88%92+4%29%289x+%2B+4%29+%3D+ax2+%E2%88%92+b

OpenStudy (anonymous):

Great!!! One more?

OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of f[g(−4)]? −12 3 −13 8

undeadknight26 (undeadknight26):

This one is pinching me.

OpenStudy (anonymous):

same

undeadknight26 (undeadknight26):

I have no idea lol

OpenStudy (anonymous):

Alright man thanks for the help!!

undeadknight26 (undeadknight26):

ANy time m8! Sorry I couldn't help all the way!